# Gradient-Based Program Repair: Fixing Bugs in Continuous Program Spaces

### v1

@@ 2 Background
Transformer Programs. In this paper, we deal with programs that can be represented both symbolically and numerically. Our symbolic program space consists of RASP (Weiss et al., 2021) programs. RASP is a programming language specifically designed for sequence processing tasks, whose primitives can be directly implemented by components of the Transformer architecture. In our experiments, our numerical program space consists of Transformer models (Vaswani et al., 2017). We leverage Tracr (Lindner et al., 2023), a compiler to translate any RASP symbolic program into an equivalent numerical program, which is a Transformer.
@@ 3.1 Compilation of Symbolic Programs to Differentiable Numerical Programs
Numerical Execution. The execution of $D_{f,\theta}$ must always be the same as the symbolic execution of the equivalent $P_{f}$, guaranteeing that program semantics are the same in both the symbolic and numerical spaces.
@@ 3.2 Gradient-Based Program Repair (GBPR)
Repair as Gradient Descent. Gradient descent acts like rolling a ball down this landscape. The initial parameters $\theta_{b}$ place the ball somewhere corresponding to the buggy program’s behavior. The gradient $\nabla_{\theta}\mathcal{L}$ points uphill towards higher loss (more incorrect behavior). By moving in the opposite direction ($-\nabla_{\theta}\mathcal{L}$), we iteratively adjust the parameters $\theta$, effectively improving the program’s behavior step-by-step towards the desired correct functionality defined by the input-output specification. Starting from the initial parameters $\theta^{(0)}=\theta_{b}$ obtained from compiling the buggy program, we iteratively update the parameters in the direction opposite to the gradient of the loss:
@@ 5 Experiments
Training and Evaluation. Each buggy transformer program is fine-tuned via supervised learning on its train split (section 4), minimizing cross-entropy correctness loss between predicted and ground-truth output sequences. We use batch size 256, learning rate $1\times 10^{-4}$, and train up to 10k epochs with early stopping (validation loss improvement $<1\times 10^{-4}$ for 10 epochs). Repaired programs are evaluated on the test set via greedy decoding (temperature 0), reporting accuracy as exact output match percentage. Experiments used multi-instance NVIDIA A100 GPUs (1/7th A100 compute, 10GB VRAM, 2 CPUs, 32GB RAM per instance/run).
Repairing Transformer Program. To evaluate the effectiveness of Gradient-Based Program Repair, we apply it to the entire RaspBugs benchmark. Our goal is to determine whether gradient-based optimization can reliably repair a wide variety of buggy transformer programs.
@@
Repair Trajectories through the Correctness Landscape. To provide further insight into how Gradient-Based Program Repair operates, we visualize repair trajectories from buggy programs to repaired ones across the numerical program space.
@@ 6 Discussion
Limitations. Our evaluation is limited to RaspBugs, our benchmark of RASP programs; broader experimentation with other programs (e.g., with Thurnherr and Scheurer (2024)) and languages is left to future work. Additionally, GBPR can only optimize parameters within the initial model architecture obtained after Tracr compilation. If repairing a bug requires changing the structure itself (e.g., adding a new attention head), our prototype could not repair the bug. Future work is needed on symbolic-to-numerical compilation to maximize the expressivity of the numerical program space.
@@ 7 Related Work
Latent Programs. Latent programs are represented in a latent space, a compressed feature space preserving meaningful data features and placing similar points adjacently. Neelakantan et al. (2015) train a Neural Programmer to recursively select operations and data sources via latent representations at each execution step. Hong et al. (2021) find that generating discrete latent codes representing high-level operations improves program synthesis accuracy when compared with token-level generation. Bonnet and Macfarlane (2024) learn a latent program space for ARC-AGI programs, and use gradient-based search to find correct programs. Liskowski et al. (2020) train an autoencoder to embed programs in a latent space, mapping them back with an evolutionary algorithm. None of these works do program repair. Beyond our focus on RASP in this paper, Gradient-Based Program Repair is conceptually applicable to other latent program representations such as the ones from this related work.
@@
Neural Surrogates. Neural surrogates (Esmaeilzadeh et al., 2012; Renda et al., 2021) are neural networks designed to approximate complex functions with better execution efficiency. Neural surrogates are typically trained on a subset of the program’s input-output space. Weber et al. (2024) train a compiler that generates neural surrogates from the program source code, bypassing the need for generating input-output examples. Any class of programs that can be fully handled by a neural surrogate can be handled by Gradient-Based Program Repair.
Learning-based Program Repair. Several works have proposed using machine learning to repair programs (Vasic et al., 2019; Chen et al., 2019). In particular, LLMs are used to repair programs both in single-turn (Xia et al., 2023; Jiang et al., 2023) and agentic (Yang et al., 2024; Wang et al., 2024) setups. Our work is different in that we focus on repairing programs in a numerical space, using gradient-based optimization in search of the correct program, rather than searching exclusively in the token space.
@@ Appendix A RaspBugs: Benchmark Details
dyck-1 Returns a sequence of ones if the sequence is balanced in regards to open and closed parenthesis, otherwise returns zeros. [(,),(,(,)] [0,0,0,0,0]

### v2

@@ 2 Background
RASP Language. RASP (Restricted Access Sequence Processing) is a domain-specific programming language for sequence processing (Weiss et al., 2021). Its key caharacteristic is that the language primitives align with the transformer-encoder model. RASP programs operate on token and position streams using selectors ($n\times n$ relations that hold when a predicate over per-position values is true), per-position aggregation (reductions along selectors), and elementwise maps. We adopt the bounded setting of RASP which assumes fixed vocabulary and a maximum sequence length.
Tracr Compiler. Tracr (Lindner et al., 2023) is an analytical, deterministic compiler from RASP to encoder-style transformer models. It programmatically constructs attention from RASP selectors (select $\to$ query/key scores high exactly where the predicate holds; values route information) and implements aggregation and maps via attentiona and MLP steps respectively. For a fixed vocabulary and sequence length, the compiled model is strictly equivalent to the RASP program ($D(\mathbf{x})=P(\mathbf{x})$ on the supported domain).
@@ 3.1 Compilation of Symbolic Programs to Differentiable Numerical Programs
Numerical Execution. The execution of $D_{f,\theta}$ must match the input-output behavior of $P_{f}$ on the supported domain, guaranteeing that program semantics agree in both the symbolic and numerical spaces.
@@ 3.2 Gradient-Based Program Repair (GBPR)
Repair as Gradient Descent. Gradient descent acts like rolling a ball down this landscape. The initial parameters $\theta_{b}$ place the ball somewhere corresponding to the buggy program’s behavior. The gradient $\nabla_{\theta}\mathcal{L}$ points uphill towards higher loss (more incorrect behavior). By moving in the opposite direction ($\nabla_{\theta}\mathcal{L}$), we iteratively adjust the parameters $\theta$, effectively improving the program’s behavior step-by-step towards the desired correct functionality defined by the input-output specification. Starting from the initial parameters $\theta^{(0)}=\theta_{b}$ obtained from compiling the buggy program, we iteratively update the parameters in the direction opposite to the gradient of the loss:
@@ 5 Experiments
5.1 Training and Evaluation.
Each buggy transformer program is fine-tuned via supervised learning on its train split (section 4), minimizing cross-entropy correctness loss between predicted and ground-truth output sequences. We use batch size 256, learning rate $1\times 10^{-4}$, and train up to 10k epochs with early stopping (validation loss improvement $1\times 10^{-4}$ for 10 epochs). Repaired programs are evaluated on the test set via greedy decoding (temperature 0), reporting accuracy as exact output match percentage. Experiments used multi-instance NVIDIA A100 GPUs (1/7th A100 compute, 10GB VRAM, 2 CPUs, 32GB RAM per instance/run).
5.2 Repairing Transformer Program.
To evaluate the effectiveness of Gradient-Based Program Repair, we apply it to the entire RaspBugs benchmark. Our goal is to determine whether gradient-based optimization can reliably repair a wide variety of buggy transformer programs.
@@
Takeaway GBPR successfully repairs buggy programs in the numerical program space. Experiments show it restores correctness for the majority of bugs across 5 out of 6 base RASP programs, achieving near-perfect repair even for initially completely broken programs.
Figure 4: Repair success rates over evaluation steps, stratified by mutation order. In GBPR, an epoch is a full forward+backward pass over the I/O samples of the training dataset; in symbolic baselines, an epoch is the symbolic evaluation of one mutated program against the same I/O samples. The idea is that all three approaches are exposed to the same amount of information per epoch, in order to drive the search. Each panel shows accuracy trajectories for GBPR, GP, and BFS (higher better). For simple bugs (orders 1-2), symbolic methods achieve higher final performance since the number of potential repairs to explore is small. For complex bugs (orders 4-5), GBPR surpasses symbolic baselines, showing how gradient-based optimization can traverse complex program spaces more efficiently than symbolic search methods.
5.3 Comparison with Symbolic Baselines.
To contextualize the effectiveness of Gradient-Based Program Repair, we compare it against two baselines that repair buggy RASP programs through search in the symbolic program space, as opposed to GBPR which operates in the numerical program space.
The first baseline employs a Genetic Programming (GP) approach with an evolutionary strategy with population size $\mu=16$ and offspring generation $\lambda=16$ per generation, operating directly on RASP program source code represented as Abstract Syntax Trees (ASTs). Starting from the buggy program, the initial population is seeded with the original bug plus random single-step mutations. Each generation proceeds as follows: (1) parent selection via tournament selection with $k=3$ (i.e., we select 3 programs at random and keep the one with the highest fitness), (2) mutation of selected parents to generate offspring according the mutation operators, and (3) replacement where the top $\mu$ individuals by fitness are retained from the combined pool of parents and offspring. The fitness function is defined as accuracy on the training dataset, computed by executing each candidate program on all input-output pairs and measuring sequence-level exact matches (ignoring the BOS token). The search terminates either when achieving 100% accuracy or after evaluating 500 programs (this corresponds to $\approx$15 generations if all of them are complete). The second baseline implements exhaustive breadth-first search (BFS) that systematically explores all possible mutations paths from the buggy program.
Critically, both baselines use the same mutation operators employed to generate the buggy programs in RaspBugs (section 4). This represents an optimistic baseline: in realistic repair scenarios, one would not have prior knowledge of the exact fault operators that introduced the bugs, whereas our baselines benefit from this oracle information.
Figure 4 shows repair success rates (test accuracy $\geq 99\%$) over epochs. In GBPR, an epoch is a full forward+backward pass over the I/O samples of the training dataset; in symbolic baselines, an epoch is the symbolic evaluation of one mutated program against the same I/O samples. The idea is that all three approaches are exposed to the same amount of information per epoch, in order to drive the search. Figure 4 is stratified by mutation order (one panel per order, the top left figure is for simple bugs based on single mutations, the bottom right is for complex bugs stacking five different mutations). Each panel plots pass rate trajectories for all three methods, revealing both asymptotic dynamics and final performance.
Asymptotic behavior. First, we see that all three methods exhibit asymptotic behavior after some time. GBPR asymptotically reaches a performance level after approximately 75 training epochs, while the two symbolic baselines continue improving through their respective evaluation budgets but at slower rates. This reflects fundamentally different search dynamics. Gradient-based optimization is able to quickly make use of the training signal via gradient, combining the information of the behavior for multiple I/O samples at once. On the contrary, symbolic methods do not have this powerful joint convergence over I/O samples, but instead are able to steadily explore the search space.
Performance by bug complexity. Examining convergence patterns across mutation orders reveals how different paradigms scale with bug complexity. For simple bugs (mutation orders 1-2), symbolic methods achieve substantially higher pass rates at convergence (top left and middle figures). BFS fixes all single-mutation bugs (order 1, top left) as it simply explores all single-step mutations from the buggy program, which is to be expected and also acts as a sanity check for our pipeline. As the number of mutations increases, the number of potential repairs grows exponentially, meaning that symbolic search becomes ever harder. Indeed, at mutation order $\geq 3$, GBPR starts to be competitive compared to symbolic search, with faster optimization but still lower final performance (top right). For complex bugs (orders 4-5), GBPR surpasses both symbolic baselines at their respective convergence points. This clearly validates the intuition that GBPR enables a joint optimization over multiple faulty locations, where a single move along the gradient in the numerical space is equivalent to multiple mutations in the symbolic space.
This pattern reflects fundamental differences in how those two search paradigms (symbolic versus numerical) scale with bug complexity. Symbolic methods are able to systematically explore the search space of small modifications, excelling at bugs requiring one ot two local changes. However, they face combinatorial challenges as the number of potential repairs grows exponentially with the number of locations to fix. Continuous optimization navigates high-dimensional parameter spaces via gradients, trading performance on local, simple bugs for the ability to handle complex, multilocation bugs via moves in the numerical space.
Takeaway While symbolic search methods (GP, BFS) excel at simple bugs, GBPR outperforms them on complex, multi-location bugs. Gradient-based optimization scales better with bug complexity by enabling joint updates across the entire program structure in the numerical program space.
@@
5.4 Repair Trajectories through the Correctness Landscape.
To provide further insight into how Gradient-Based Program Repair operates, we visualize repair trajectories from buggy programs to repaired ones across the numerical program space.
@@
Takeaway Repair trajectory visualizations confirm that GBPR meaningfully navigates the numerical program space. By following the gradient of the correctness loss, the optimization process steadily moves the program from buggy to correct behavior, with intermediate steps being progressively closer to specification.
@@ 6 Discussion
Limitations. Our evaluation is limited to RaspBugs, our benchmark of RASP programs; broader experimentation with other programs (e.g., with Thurnherr and Scheurer (2024)) and languages is left to future work. Additionally, GBPR can only optimize parameters within the initial model architecture obtained after Tracr compilation (see scope of Tracr in section 2). If repairing a bug requires changing the structure itself (e.g., adding a new attention head), our prototype could not repair the bug. Future work is needed on symbolic-to-numerical compilation to maximize the expressivity of the numerical program space. A further practical limitation is the reliance on large input-output datasets to define the correctness loss and drive repair (we use up to 50k pairs per program; see section 5). Real-world test suites may be smaller and incomplete, limiting applicability and increasing overfitting risk under sparse specifications.
@@ 7 Related Work
Latent Programs. Latent programs are represented in a latent space, a compressed feature space preserving meaningful data features and placing similar points adjacently. Neelakantan et al. (2015) train a Neural Programmer to recursively select operations and data sources via latent representations at each execution step. Hong et al. (2021) find that generating discrete latent codes representing high-level operations improves program synthesis accuracy when compared with token-level generation. Bonnet and Macfarlane (2024) learn a latent program space for ARC-AGI programs, and use gradient-based search to find correct programs. Liskowski et al. (2020) train an autoencoder to embed programs in a latent space, mapping them back with an evolutionary algorithm. None of these works do program repair. Beyond our focus on RASP in this paper, Gradient-Based Program Repair is conceptually applicable to other latent program representations such as the ones from this related work. Lastly, Balog et al. (2020) introduce a neural synthesizer with a differentiable fixer that iteratively revises symbolic programs using latent program representation intermediates; unlike GBPR, their method relies on the discretization of each edit step rather than optimizing the program fully in a continuous space.
@@
Symbolic vs Numerical Program Spaces. The mapping between symbolic and numerical program spaces is a key component in gradient-based program repair, and a topic of earlier research. Neural surrogates (Esmaeilzadeh et al., 2012; Renda et al., 2021) are neural networks designed to approximate complex programs and are typically trained on a subset of the input-output space. Weber et al. (2024) learn to compile source programs directly into neural surrogates, bypassing the need for generating input-output examples. Smooth interpretation (Chaudhuri and Solar-Lezama, 2010) takes a different route: it smooths the semantics of a symbolic program to expose continuous parameters amenable to gradient-based optimization. Other works focus on mapping from numerical to symbolic program spaces. Cranmer et al. (2020) propose symbolic regression for extracting explicit symbolic models from learned models by relying on inductive biases matching to nature of the problem domain. Thurnherr and Riesen (2024); Langosco et al. (2024); Upadhyay et al. (2025) study the decompilation of Transformer models into symbolic programs by learning meta-decompiler models. Given such as mapping, the core novelty of GBPR is to explore the numerical program space for finding patches, using gradient descent.
Learning-based Program Repair. Several works have proposed using machine learning to repair programs (Long and Rinard, 2016; Vasic et al., 2019; Chen et al., 2019). In particular, LLMs are used to repair programs both in single-turn (Xia et al., 2023; Jiang et al., 2023) and agentic (Yang et al., 2024; Wang et al., 2024) setups. Our work is different in that we focus on repairing programs in a numerical space, using gradient-based optimization in search of the correct program, rather than searching exclusively in the token space.
@@ Appendix A RaspBugs: Benchmark Details
dyck-1 Returns a sequence of ones if the sequence is balanced in regards to open and closed parenthesis, otherwise returns zeros. [(,),$(,$)] [0,0,0,0,0]